What is the amplitude, period, phase shift, and vertical shift of:
y=3sin4(x-(π/2))+2
I am completely stumped on this and my solutions aren't making sense. Any help will be greatly appreciated.

Answer by lwsshak3(11628)

What is the amplitude, period, phase shift, and vertical shift of:
y=3sin4(x-(π/2))+2
..
Standard form for the sine function: y=Asin(Bx-C), with A=amplitude,2π/B=Period, C/B=phase-shift.
For given sine function:
y=3sin4(x-(π/2))+2
rewrite so it looks more like the standard form by expanding 4(x-(π/2))
y=3sin(4x-2π)+2
A=3=amplitude
B=4
period=2π/B=2π/4=π/2
C=2π
phase-shift=C/B=2π/4=π/2 (to the right)
note:( in the given sine function, 4 is factored out to show the phase-shift of π/2)
+2=vertical shift of the curve up 2 units